# Harnessing the Power of Neural Operators with Automatically Encoded Conservation Laws

### v2

@@ Appendix A Related concepts of differential forms
Let $k$ be a non-negative integer and $\mu(\mathbf{x})=f(x_{1},\cdots,x_{p})$ be a smooth function on $\mathbb{R}^{p}$, a monomial $k-$form on $\mathbb{R}^{p}$ is an expression $\mu\,dx_{i_{1}}\wedge\cdots\wedge dx_{i_{k}}=\mu dx_{I}$, where $I:=\{i_{1},\cdots,i_{k}$. A $k-$form is a sum of monomial $k-$forms, which will be denoted as
@@
The Hodge normal-$\star$-operator on $\mathbb{R}^{p}$ is a function that takes a $k-$form to a $(p-k)-$form defined as follows:
@@
Given a vector-valued function on $\mathbb{R}^{p}$ $\mathbf{u}(\mathbf{x})=[u_{1}(\mathbf{x}),\cdots,u_{p}(\mathbf{x})]$, its divergence $\text{div}(\mathbf{u})=\sum_{i=1}^{p}\frac{\partial{u_{i}}}{\partial x_{i}}$ can be defined via exterior derivative:
@@
Appendix B Data generation and training strategies
B.1 Example 1 – incompressible Navier-Stokes equation
For the two case study datasets in the incompressible Navier-Stokes example, we generate a total of 1,200 samples using the pseudo-spectral Crank-Nicolson solver available in Li et al. (2020c). We then split the generated dataset into 1,000, 100 and 100 for training, validation and testing, respectively. A histogram demonstration of the dataset distributions in small, medium and large data regimes is illustrated in Figure 3, where the test set of the small dataset of ntrain=10 samples exhibits a wider data distribution compared to its training dataset, which is aimed to test the data efficiency and out-of-distribution performances of the learned models. The fluid viscosity employed in the physics solver is $\nu=10^{-4}$, and the timestep size is $\Delta t=10^{-4}$ s. The solutions are obtained on a $256\times 256$ spatial grid and the total duration of the simulation is 24 s. The obtained solutions are then downsampled to a $64\times 64\times 30$ grid, with the 3rd dimension being the temporal dimension. Note that, in downsampling the spatial dimensions, we employ a $2\times 2$ mean pooling. This strategy is suggested in Helwig et al. (2023) to mitigate the spurious numerical errors not existed in the original data.
@@ B.1 Example 1 – incompressible Navier-Stokes equation
Remark: The complexity of our clawNOs is very simular to their NO counterparts. Taking clawFNOs for example, the trainable part of the clawNO model consists of two fixed-size MLPs for the lifting layer and the projection layer, and $L$ numbers of Fourier layers between them. Denote $d_{u}$ as the input function dimension, $H$ as the latent dimension after lifting, $M$ as the total number of grids, $m$ as the number of Fourier modes on each dimension after truncation (which is often taken as half of the number of grids in each dimension, $M^{1/p}$, in practice), and $p$ as the problem dimension. During the lifting layer a vector valued input function taking values in $\mathbb{R}^{d_{u}}$ is linearly and locally mapped to the first layer feature function $\mathbf{h}(\cdot,0)$ taking values in $\mathbb{R}^{H}$, and hence the number of trainable parameters is $Hd_{u}+H$. Then, each iterative Fourier layers involves the integral with a trainable kernel weight matrix in the Fourier domain, which is of size $2H^{2}m^{p}=2^{1-p}H^{2}M$, and a local linear transformation which involves $H^{2}+H$ numbers of trainable parameters. Then, the last iterative layer feature function, $\mathbf{h}(\cdot,L)$, is projected to the skew symmetric matrix-valued function $\mu$ with a two-layer MLP. Since the skew symmetric matrix-valued function $\mu$ is of degree of freedom $p(p-1)/2$ at each point $\mathbf{x}$, the projection layer maps a size $H$ input vector to a size $p(p-1)/2$ vector. Assume that the hidden layer of this MLP is of $d_{Q}$ neurons, the total number of trainable parameters in projection layer will be $Hd_{Q}+d_{Q}+d_{Q}p(p-1)/2+p(p-1)/2$. Finally, $\mu$ will go through the pre-calculated differentiation layer, with $p^{2}(p-1)M^{2}/2=p^{2}(p-1)m^{2p}/2$ numbers of non-trainable parameters in the FNO case. From the above calculation, we can see that clawFNO involves $(d_{Q}+H(d_{u}+d_{Q}+L+1)+LH^{2})+\dfrac{(d_{Q}+1)}{2}p(p-1)+2LH^{2}m^{p}$ numbers of trainable parameters, while the vanilla FNO involves $(d_{Q}+H(d_{u}+d_{Q}+L+1)+LH^{2})+(d_{Q}+1)p+2LH^{2}m^{p}$ numbers of trainable parameters. Therefore, the number of trainable parameters in clawFNO and FNO only differs in the second part of their projection layer, where clawFNO has $\dfrac{(d_{Q}+1)}{2}p(p-1)$ numbers of parameters and FNO has $(d_{Q}+1)p$. When $p>3$, clawFNO will have a larger number of trainable parameters. However, we want to point out that since the number of parameter in the iterative layer ($2LH^{2}m^{p}$) grows exponentially with dimension $p$, it dominates the cost, and the differences between clawFNO and FNO are negligible. This is consistent with what we observed in Table 5: the number of trainable parameters and the GPU cost of clawNOs and their counterparts are almost the same. During the inference, the non-trainable parameters in clawNO will play a role and we therefore observe an increase in the inference time.
@@
B.2 Example 2 – radial dam break
In generating the radial dam break dataset, we closely follow the numerical procedure in Takamoto et al. (2022), where we slightly modified the code to output the velocity fields in addition to the water height. A total of 1,000 samples are generated and are subsequently split into $n_{train}$, 100, 100 and training, validation, and testing, respectively, with $n_{train}$ the size of the training dataset depending on the adopted data regime. We run the numerical simulation on a $128\times 128\times 100$ grid and downsample the obtained solution to $32\times 32\times 25$ for training, where the first two dimensions are the spatial dimensions and the last is the temporal dimension. Analogous to the incompressible NS dataset, we perform $2\times 2$ mean pooling in downsampling the spatial dimensions to maintain symmetry in data.
B.3 Example 3 – atmospheric modeling
@@
For numerical stability, an implicitly calculated horizontal diffusion term of the form $-\nu\nabla^{8}\zeta,-\nu\nabla^{8}\mathcal{D}$, is added to the vorticity and the divergence equation, respectively. The power-4 Laplacian is chosen to be very scale-selective such that energy is only removed at the highest wave numbers, keeping the simulated flow otherwise largely unaffected. In practice, we use a non-dimensional Laplace operator ~$\tilde{\nabla}^{2}=R^{2}\nabla^{2}$, such that the diffusion coefficient becomes an inverse time scale of $\tau=2.4~{}\text{hours}$. The shallow water equations, equation 12, do not have a forcing or drag term such that the horizontal diffusion is the only term through which the system loses energy over time. The shallow water equations are otherwise energy conserving.
@@
B.4 Example 4 – constitutive modeling of material deformation
We generate the material deformation dataset as follows. Firstly, we generate a solution $A$ from a steady-state 2D Darcy flow with the diffusion coefficient specified as a random Gaussian field, following the numerical procedure in Li et al. (2020c). The ground truth $\mathbf{u}$ is generated as
$\mathbf{u}=(\dfrac{\partial A}{\partial y}\text{ ,}-\dfrac{\partial A}{\partial x})\text{ ,}$
which is guaranteed divergence-free, and the body load $\mathbf{f}$ is then computed following equation 13. The derivatives are numerically computed using FFT on a regular mesh. A total of 300 samples are generated and are subsequently split into $n_{train}$, 100, 100 for training, validation, and testing, respectively, with $n_{train}$ the size of the training dataset depending on the adopted data regime. We run the numerical simulation on a 80$\times$80 uniform mesh and downsample the solution to a grid of 246 spatial points through interpolation on a circular domain of radius $0.4$ and centers at the origin point.
B.5 Training strategies
@@
Appendix C Rollout visualizations
We illustrate the rollout of randomly selected trajectories in the test dataset using clawNO predictions, along with the comparisons against ground truth data and the corresponding absolute errors. The rollouts of incompressible NS, radial dam break, and atmospheric modeling are showcased in Figure 6, Figures 8 and 9, Figures 11 and 2, respectively. We also showcase the material deformation prediction in 6 different test samples in Figure 15. To provide a visual comparison across models, we plot the final-step prediction of all the models in the incompressible NS example in Figure 7, the radial dam break example in Figure 10, and the atmospheric modeling example in Figure 14.
@@ Appendix C Rollout visualizations
Appendix D Detailed derivations
D.1 Proof of theorem 3.1
Since the Fourier spectral differentiation error estimate is a direct result of Trefethen (2000, Page 34), we provide the detailed derivation of equation 7 in this section. According to the basic properties of Fourier transform, if $f$ is a differential and periodic function on $[0,L]$ with its Fourier representation:
@@ Appendix D Detailed derivations
As such, equation 7 can be obtained by applying the above property to approximate every derivative term $\dfrac{\partial\mu^{(N^{(1)},\cdots,N^{(p)})}_{jk}}{\partial x_{k}}$.
D.2 Proof of theorem 3.2

### v3

@@ Appendix A Related concepts of differential forms
Let $k$ be a non-negative integer and $\mu(\mathbf{x})=f(x_{1},\cdots,x_{p})$ be a smooth function on p, a monomial $k-$form on p is an expression $\mu\,dx_{i_{1}}\wedge\cdots\wedge dx_{i_{k}}=\mu dx_{I}$, where $I:=\{i_{1},\cdots,i_{k}$. A $k-$form is a sum of monomial $k-$forms, which will be denoted as
@@
The Hodge $\star$-operator on p is a function that takes a $k-$form to a $(p-k)-$form defined as follows:
@@
Given a vector-valued function on p $\mathbf{u}(\mathbf{x})=[u_{1}(\mathbf{x}),\cdots,u_{p}(\mathbf{x})]$, its divergence $\text{div}(\mathbf{u})=\sum_{i=1}^{p}\frac{\partial{u_{i}}}{\partial x_{i}}$ can be defined via exterior derivative:
@@
Appendix B Equivariance of the differentiation layer
The projection layer ($\mu$ to $\mathbf{u}$) is realized as a numerical divergence operator. Therefore, under the translation of coordinates, $\tilde{\mathbf{x}}=\mathbf{x}+g$, we have
$\nabla_{\tilde{\mathbf{x}}}\cdot\mu_{j}=\sum_{k}\dfrac{\partial}{\partial\tilde{x}_{k}}\mu_{jk}=\sum_{k}\dfrac{\partial}{\partial{x}_{k}}\mu_{jk}=\nabla_{{\mathbf{x}}}\cdot\mu_{j}.$
That means, the divergence operator is invariant under coordinate translation, and therefore $\mathcal{D}$ is also invariant. On the other hand, under the rotation of coordinates, $\tilde{\mathbf{x}}=R\mathbf{x}$, we have the chain rule: $\dfrac{\partial}{\partial{x}_{k}}=\sum_{l}R_{lk}\dfrac{\partial}{\partial\tilde{x}_{l}}$. Therefore,
$\nabla_{\tilde{\mathbf{x}}}\cdot(R\mu_{j})=\sum_{k}\dfrac{\partial}{\partial\tilde{x}_{k}}(R\mu_{j})_{k}=\sum_{k}\dfrac{\partial}{\partial\tilde{x}_{k}}\sum_{l}R_{kl}\mu_{jl}=\sum_{l}\dfrac{\partial}{\partial{x}_{l}}\mu_{jl}=\nabla_{{\mathbf{x}}}\cdot\mu_{j}.$
That means, the coefficients in $\mathcal{D}$ is also invariant under coordinate rotation.
When the previous layers are translational invariant and rotational equivariant, we have $\tilde{\mu}_{j}(\tilde{\mathbf{x}})=\tilde{\mu}_{j}(R\mathbf{x}+g)=R\mu_{j}(\mathbf{x})$ for the $j-$th row of $\mu$. Therefore, $\mathcal{D}[\tilde{\mu}](\tilde{\mathbf{x}})=R\mathcal{D}[\mu](\mathbf{x})$, which results in the translational invariant and rotational equivariant output.
Appendix C Data generation and training strategies
C.1 Example 1 – incompressible Navier-Stokes equation
For the two case study datasets in the incompressible Navier-Stokes example, we generate a total of 1,200 samples using the pseudo-spectral Crank-Nicolson solver available in Li et al. (2020c). Here, the models are trained to predict the two velocity components in the two directions. We then split the generated dataset into 1,000, 100 and 100 for training, validation and testing, respectively. A histogram demonstration of the dataset distributions in small, medium and large data regimes is illustrated in Figure 3, where the test set of the small dataset of ntrain=10 samples exhibits a wider data distribution compared to its training dataset, which is aimed to test the data efficiency and out-of-distribution performances of the learned models. The fluid viscosity employed in the physics solver is $\nu=10^{-4}$, and the timestep size is $\Delta t=10^{-4}$ s. The solutions are obtained on a $256\times 256$ spatial grid and the total duration of the simulation is 24 s. The obtained solutions are then downsampled to a $64\times 64\times 30$ grid, with the 3rd dimension being the temporal dimension. Note that, in downsampling the spatial dimensions, we employ a $2\times 2$ mean pooling. This strategy is suggested in Helwig et al. (2023) to mitigate the spurious numerical errors not existed in the original data.
@@ C.1 Example 1 – incompressible Navier-Stokes equation
Remark: The complexity of our clawNOs is very simular to their NO counterparts. Taking clawFNOs for example, the trainable part of the clawNO model consists of two fixed-size MLPs for the lifting layer and the projection layer, and $L$ numbers of Fourier layers between them. Denote $d_{u}$ as the input function dimension, $H$ as the latent dimension after lifting, $M$ as the total number of grids, $m$ as the number of Fourier modes on each dimension after truncation (which is often taken as half of the number of grids in each dimension, $M^{1/p}$, in practice), and $p$ as the problem dimension. During the lifting layer a vector valued input function taking values in ${}^{d_{u}}$ is linearly and locally mapped to the first layer feature function $\mathbf{h}(\cdot,0)$ taking values in H, and hence the number of trainable parameters is $Hd_{u}+H$. Then, each iterative Fourier layers involves the integral with a trainable kernel weight matrix in the Fourier domain, which is of size $2H^{2}m^{p}=2^{1-p}H^{2}M$, and a local linear transformation which involves $H^{2}+H$ numbers of trainable parameters. Then, the last iterative layer feature function, $\mathbf{h}(\cdot,L)$, is projected to the skew symmetric matrix-valued function $\mu$ with a two-layer MLP. Since the skew symmetric matrix-valued function $\mu$ is of degree of freedom $p(p-1)/2$ at each point $\mathbf{x}$, the projection layer maps a size $H$ input vector to a size $p(p-1)/2$ vector. Assume that the hidden layer of this MLP is of $d_{Q}$ neurons, the total number of trainable parameters in projection layer will be $Hd_{Q}+d_{Q}+d_{Q}p(p-1)/2+p(p-1)/2$. Finally, $\mu$ will go through the pre-calculated differentiation layer, with $p^{2}(p-1)M^{2}/2=p^{2}(p-1)m^{2p}/2$ numbers of non-trainable parameters in the FNO case. From the above calculation, we can see that clawFNO involves $(d_{Q}+H(d_{u}+d_{Q}+L+1)+LH^{2})+\dfrac{(d_{Q}+1)}{2}p(p-1)+2LH^{2}m^{p}$ numbers of trainable parameters, while the vanilla FNO involves $(d_{Q}+H(d_{u}+d_{Q}+L+1)+LH^{2})+(d_{Q}+1)p+2LH^{2}m^{p}$ numbers of trainable parameters. Therefore, the number of trainable parameters in clawFNO and FNO only differs in the second part of their projection layer, where clawFNO has $\dfrac{(d_{Q}+1)}{2}p(p-1)$ numbers of parameters and FNO has $(d_{Q}+1)p$. When $p>3$, clawFNO will have a larger number of trainable parameters. However, we want to point out that since the number of parameter in the iterative layer ($2LH^{2}m^{p}$) grows exponentially with dimension $p$, it dominates the cost, and the differences between clawFNO and FNO are negligible. This is consistent with what we observed in Table 5: the number of trainable parameters and the GPU cost of clawNOs and their counterparts are almost the same. During the inference, the non-trainable parameters in clawNO will play a role and we therefore observe an increase in the inference time.
@@
C.2 Example 2 – radial dam break
In generating the radial dam break dataset, we closely follow the numerical procedure in Takamoto et al. (2022), where we slightly modify the code to output the velocity fields in addition to the water height. The model then aims to learn three channels of $(h,hu_{x},hu_{y})$. A total of 1,000 samples are generated and are subsequently split into $n_{train}$, 100, 100 and training, validation, and testing, respectively, with $n_{train}$ the size of the training dataset depending on the adopted data regime. We run the numerical simulation on a $128\times 128\times 100$ grid and downsample the obtained solution to $32\times 32\times 25$ for training, where the first two dimensions are the spatial dimensions and the last is the temporal dimension. Analogous to the incompressible NS dataset, we perform $2\times 2$ mean pooling in downsampling the spatial dimensions to maintain symmetry in data. In addition, we demonstrate the zero-shot super-resolution predictability of clawNO in Figure 5, where we use the clawFNO trained on $32\times 32$ spatial resolutions to directly make predictions on $128\times 128$ grids.
Figure 5: Demonstration of zero-shot super resolution of clawFNO using the radial dam break dataset, where clawFNO is trained on $32\times 32$ spatial resolutions and directly make predictions on $128\times 128$ grids.
C.3 Example 3 – atmospheric modeling
@@
For numerical stability, an implicitly calculated horizontal diffusion term of the form $-\nu\nabla^{8}\zeta,-\nu\nabla^{8}\mathcal{D}$, is added to the vorticity and the divergence equation, respectively. The power-4 Laplacian is chosen to be very scale-selective such that energy is only removed at the highest wave numbers, keeping the simulated flow otherwise largely unaffected. In practice, we use a non-dimensional Laplace operator ~$\tilde{\nabla}^{2}=R^{2}\nabla^{2}$, such that the diffusion coefficient becomes an inverse time scale of $\tau=2.4~{}\text{hours}$. The shallow water equations, (4.3), do not have a forcing or drag term such that the horizontal diffusion is the only term through which the system loses energy over time. The shallow water equations are otherwise energy conserving.
@@
In this context, the mass conservation in spherical coordinates can be expressed as: $\frac{\partial\eta}{\partial t}+\nabla\cdot(\mathbf{u}h)=0$, where $\eta$ is the displacement from the atmosphere’s rest height $H=8500~{}\text{m}$, $h=\eta+H-H_{o}$ is the layer thickness, $H_{o}=H_{o}(\lambda,\phi)$ is the Earth’s orography, and $\lambda$, $\phi$ are the longitude and colatitude, respectively. We note that $\frac{\partial\eta}{\partial t}=\frac{\partial h}{\partial t}$, and $\nabla\cdot(\mathbf{u}h)$ in the spherical coordinate can be rewritten as: $\nabla\cdot(\mathbf{u}h)=\frac{1}{R\sin\theta}\frac{\partial(u_{x}h)}{\partial\lambda}+\frac{1}{R\sin\theta}\frac{\partial(u_{y}h\sin\theta)}{\partial\theta}$ (here definitions using latitudes replace the $\sin$ with a $\cos$). Based on this formulation, we first rescale the equation by $R$, then set the clawNO model to learn $(u_{x}h,u_{y}h\sin\theta,Rh\sin\theta)$ and require the output to be divergence-free:
$(\partial_{\lambda},\partial_{\theta},\partial_{t})\cdot(u_{x}h,u_{y}h\sin\theta,Rh\sin\theta)=\dfrac{\partial(u_{x}h)}{\partial\lambda}+\dfrac{\partial(u_{y}h\sin\theta)}{\partial\theta}+\dfrac{\partial(Rh\sin\theta)}{\partial t}=\dfrac{\partial(u_{x}h)}{\partial\lambda}+\dfrac{\partial(u_{y}h\sin\theta)}{\partial\theta}+R\sin\theta\dfrac{\partial\eta}{\partial t}=0.$
As such, the mass conservation equation is guaranteed. Once the output field $\mathbf{u}=(u_{x}h,u_{y}h\sin\theta,Rh\sin\theta)$ is obtained, we post-process the prediction to obtain $(u_{x},u_{y},h)$ (e.g., by dividing the second component $u_{y}h\sin\theta$ by the third component $Rh\sin\theta$ and multiplying the result by $R$, to obtain $u_{y}$). In this way, we can leave our skew symmetric matrix and numerical differentiation layer unchanged while being able to handle spherical data.
C.4 Example 4 – constitutive modeling of material deformation
We generate the material deformation dataset using an incompressible Mooney–Rivlin material model. We run the numerical simulation on a 80$\times$80 uniform mesh and downsample the solution to a grid of 246 spatial points through interpolation on a circular domain of radius $0.4$ that centers at the origin (cf. Figure 7). A total of 300 samples are generated and are subsequently split into $n_{train}$, 100, 100 for training, validation, and testing, respectively, with $n_{train}$ the size of the training dataset depending on the adopted data regime.
In this example, we impose the kinematic condition of incompressibility with infinitesimal deformation, $div(\mathbf{u})=0$. This is equivalent to the mass conservation law only under the assumption that the density remains a constant, i.e., $\rho=const$, and then $div(\mathbf{u})=0$ guarantees that the total mass of a volume does not change under any applied deformation, i.e., its total mass is conserved. In solid mechanics, the deformation gradient writes $\mathbf{F}=\mathbf{I}+\nabla\mathbf{u}$, and the total mass of a volume does not change when $det(\mathbf{F})\equiv 1$. Under an infinitesimal deformation assumption, one has $\dfrac{\partial u_{i}}{\partial x_{j}}\dfrac{\partial u_{k}}{\partial x_{l}}\approx 0$, and therefore $det(\mathbf{F})\approx 1+div(\mathbf{u})$. That means, the incompressibility condition is equivalent to $div\mathbf{u}=0$.
Figure 7: An illustration of the grid points and circular domain uused in example 4.
C.5 Training strategies
@@
Appendix D Rollout visualizations
We illustrate the rollout of randomly selected trajectories in the test dataset using clawNO predictions, along with the comparisons against ground truth data and the corresponding absolute errors. The rollouts of incompressible NS, radial dam break, and atmospheric modeling are showcased in Figure 8, Figures 10 and 11, Figures 13 and 2, respectively. We also showcase the material deformation prediction in 6 different test samples in Figure 17. To provide a visual comparison across models, we plot the final-step prediction of all the models in the incompressible NS example in Figure 9, the radial dam break example in Figure 12, and the atmospheric modeling example in Figure 16.
@@ Appendix D Rollout visualizations
Appendix E Detailed error estimates for the differentiation layer
E.1 Proof of theorem 3.1
Since the Fourier spectral differentiation error estimate is a direct result of Trefethen (2000, Page 34), we provide the detailed derivation of (7) in this section. According to the basic properties of Fourier transform, if $f$ is a differential and periodic function on $[0,L]$ with its Fourier representation:
@@ Appendix E Detailed error estimates for the differentiation layer
As such, (7) can be obtained by applying the above property to approximate every derivative term $\dfrac{\partial\mu^{(N^{(1)},\cdots,N^{(p)})}_{jk}}{\partial x_{k}}$.
E.2 Proof of theorem 3.2
@@
E.3 Performance comparison of different numerical differentiation schemes
To evaluate the empirical performance of the differentiation layer, we compare the performance of the FFT and meshfree methods discussed in Theorems 3.1 and 3.2. In particular, we take an analytical divergence-free function, $\mathbf{f}=[\sin x\sin y,\cos x\cos y]$, and use the numerical differentiation layer to evaluate its divergence and report the $L^{2}$ error. This result can be found in Figure 18. Here, the $L^{2}$ divergence errors using FFT remain near machine precision irregardless of the variation of grid sizes, because the spectrum generation is exact on the provided analytical function. On the other hand, for the meshfree method we have generated quadrature rules for $5-$th order polynomials, i.e., $m=5$. A $6-$th order convergence is observed, which is consistent with the analysis in Theorem 3.2.
Figure 18: Numerical error of the differentiation layer, with refinement on grids.